### v5

# Kernelized Concept Erasure

Abstract
The representation space of neural models for textual data emerges in an unsupervised manner during training. Understanding how those representations encode human-interpretable concepts is a fundamental problem. One prominent approach for the identification of concepts in neural representations is searching for a linear subspace whose erasure prevents the prediction of the concept from the representations. However, while many linear erasure algorithms are tractable and interpretable, neural networks do not necessarily represent concepts in a linear manner. To identify non-linearly encoded concepts, we propose a kernelization of a linear minimax game for concept erasure. We demonstrate that it is possible to prevent specific non-linear adversaries from predicting the concept. However, the protection does not transfer to different non-linear adversaries. Therefore, exhaustively erasing a non-linearly encoded concept remains an open problem.
https://github.com/shauli-ravfogel/adv-kernel-removal
1 Introduction
Large neural networks in NLP produce real-valued representations that encode the bit of human language that they were trained on, e.g., words, sentences, or text grounded in images. For instance, GloVe (Pennington et al., 2014) produces real-valued representations of isolated words, BERT (Devlin et al., 2019) produces real-valued representations of sentences, and VilBERT produces real-valued representations of visually grounded language (Lu et al., 2019; Bugliarello et al., 2021). These real-valued representations naturally encode various properties of the objects they represent. For instance, a good representation of the first author’s laptop computer ought to encode its manufacturer, size, and color somewhere among its real values.
We now describe the premise of our paper in more detail. We adopt the notion of a concept due to Gärdenfors (2000). For Gärdenfors, objects can be thought of as having a geometric representation. Different dimensions in the representation space that objects inhabit might correspond to their color, size, and shape. Gärdenfors then goes further and defines a concept as a convex region of the representation space. [^1] In this work, we exclusively focus on linear subspaces, which are convex, rather than more general convex regions. Building on Gärdenfors’ notion of a concept, this paper studies a task that we refer to as concept erasure.
[…]
Concept erasure is tightly related to concept identification. Once we have successfully removed a given concept, e.g., color, from a representation $\mathbf{x}_{n}$, it is reasonable to argue that the erasure function $r$ has meaningfully identified the concept within the representation space. In the rest of the paper, we will say that $r(\cdot)$ neutralizes the concept in the representation space. For instance, we say that the $r$ in our example neutralizes the concept of a laptop’s color. It follows that concept identification is related to bias mitigation Bolukbasi et al. (2016); Gonen and Goldberg (2019); Maudslay et al. (2019), e.g., one may want to identify and remove the gender encoded in learned word representations produced by an embedding method such as word2vec (Mikolov et al., 2013) or GloVe (Pennington et al., 2014). Indeed, the empirical portion of this paper will focus on removing gender from word representations in order to mitigate bias.
Previous work on concept erasure (Ravfogel et al., 2021) focuses on the linear case, i.e., where $r$ is a linear function. While linear concept erasure methods have certainly found success Bolukbasi et al. (2016), there is no a-priori reason to suspect that neural networks encode concepts in a linear manner. In this work, we take the first step toward the goal of identifying a non-linear function $r(\cdot)$ and a corresponding non-linearly encoded concept subspace. We directly build on Ravfogel et al. (2022), who cast linear concept erasure as a minimax game. Under their formulation, the function $r(\cdot)$ learns to remove the concept, while an adversary tries to predict the concept. We extend their work by deriving a class of general minimax games based on kernelization that largely maintains the tractability of the linear approach. Our kernelized method performs concept erasure in a reproducing kernel Hilbert space, which may have a much higher dimensionality (Schölkopf and Smola, 2002) and correspond to a non-linear subspace of the original representation space.
Empirically, we experiment with gender erasure from GloVe and BERT representations. We show that a kernelized adversary can classify the gender of the representations with over 99% accuracy if $r(\cdot)$ is taken to be a linear function. This gives us concrete evidence that gender is indeed encoded non-linearly in the representations. We further find that solving our kernelized minimax game yields an erasure function $r(\cdot)$ that protects against an adversary that shares the same kernel. However, we also find that it is difficult to protect against all kernelized adversaries at once: Information removed by one kernel type can be recovered by adversaries using other kernel types. That is, the gender concept is not exclusively encoded in a space that corresponds to any one kernel. This suggests that non-linear concept erasure is very much an open problem.
2 Linear Concept Erasure
We provide an overview of the linear minimax formulation before we introduce its kernelization.
2.1 Notation
Let ${\mathcal{D}}=\{(y_{n},\mathbf{x}_{n})\}_{n=1}^{N}$ be a dataset of $N$ concept–representation pairs, where the labels $y_{n}$ represent the concept to be neutralized. The goal of linear concept erasure is to learn a linear erasure function $r(\cdot)$ from ${\mathcal{D}}$ such that it is impossible to predict $y_{n}$ from the modified representations $r(\mathbf{x}_{n})$. We focus on classification, where the concept labels $y_{n}$ are derived from a finite set $\{1,\dots,V\}$ of $V$ discrete values, and the representations $\mathbf{x}_{n}\in\mathbb{R}^{D}$ are $D$-dimensional real column vectors. To predict the concepts labels $y_{n}$ from the representations $\mathbf{x}_{n}$, we make use of (and later kernelize) classifiers that are linear models, i.e., classifiers of the form $\widehat{y}_{n}={\boldsymbol{\theta}}^{\top}\mathbf{x}_{n}$ where ${\boldsymbol{\theta}}\in\Theta\subseteq\mathbb{R}^{D}$ is a column vector of parameters that lives in a compact set $\Theta$. We also consider an arbitrary loss function $\ell(\cdot,\cdot)\geq 0$ where $\ell(y_{n},\widehat{y}_{n})$ tells us how close the prediction $\widehat{y}_{n}$ is to $y_{n}$.
Using this notation, linear concept erasure is realized by identifying a linear concept subspace whose neutralization, achieved through an orthogonal projection matrix, prevents the classifier from predicting the concept. We define $\mathcal{P}_{k}$ as the set of all $D\times D$ orthogonal projection matrices that neutralize a rank $k$ subspace. More formally, we have that $P\in\mathcal{P}_{k}\leftrightarrow P=I_{D}-W^{\top}W,W\in\mathbb{R}^{k\times D},WW^{\top}=I_{k}$, where $I_{k}$ denotes the $k\times k$ identity matrix and $I_{D}$ denotes the $D\times D$ identity matrix. We say that the matrix $P$ neutralizes the $k$-dimensional rowspace of $W$.
2.2 A Linear Minimax Game
Following this formalization, it is natural to define a minimax game (Neumann and Morgenstern, 1944) between a projection matrix $P\in\mathcal{P}_{k}$ that aims to remove the concept subspace, and a linear model parameterized by ${\boldsymbol{\theta}}$ that aims to recover it:
$\min_{{\boldsymbol{\theta}}\in\Theta}\,\max_{{P\in\mathcal{P}_{k}}}\,\sum_{n=1}^{N}\ell\Big{(}y_{n},\,{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}\Big{)}$ (1)
This is a special case of the general adversarial framework (Goodfellow et al., 2014). However, in our case, the predictor ${\boldsymbol{\theta}}$ does not interact with the original input $\mathbf{x}_{n}$. Instead, the classifier attempts to predict the concept label $y_{n}$ from $P\mathbf{x}_{n}$. We now give a concrete example of the linear concept: When we have a binary logistic loss ($V=2$) and a $k$-dimensional neutralized subspace, the game takes the following form:
$\min_{{\boldsymbol{\theta}}\in\mathbb{R}^{D}}$ $\max_{P\in\mathcal{P}_{k}}\,\,-\Bigg{(}\sum_{n=1}^{N}y_{n}\log\frac{\exp{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}}{1+\exp{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}}$ (2)
$+\Big{(}1-y_{n}\Big{)}\log\frac{1}{1+\exp{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}}\Bigg{)}$
In this paper, we focus on $k=1$, i.e., we aim to identify a 1-dimensional concept subspace. [^2] This construction can be extended to subspaces of dimension $k>1$ using a version of the Gram–Schmidt process in the RKHS. The optimization over $\mathcal{P}_{k}$ in Eq. 1 renders the game non-convex. To address this issue, Ravfogel et al. (2022) propose a concave relaxation of the objective:
$\min_{{\boldsymbol{\theta}}\in\Theta}\,\max_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}P\in\mathcal{F}_{k}}}\,\sum_{n=1}^{N}\ell\Big{(}y_{n},\,{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}\Big{)}$ (3)
where the relaxation is shown in gray. In words, instead of optimizing over rank-$k$ projection matrices, a non-convex set, we optimize over its convex hull, the Fantope (Boyd and Vandenberghe, 2014):
$\mathcal{F}_{k}=\{A\in\mathcal{S}^{D}\mid I_{D}\succcurlyeq A\succcurlyeq 0,\mathrm{tr}(A)=k\}$ (4)
3 Non-linear Concept Erasure
It has often been shown that human-interpretable concepts, and in particular gender, are encoded non-linearly in the representation space (Gonen and Goldberg, 2019; Ravfogel et al., 2020). However, prior work on concept erasure, e.g., the method discussed in § 2, assumes that concepts are encoded linearly. Our goal is to extend the game defined in Eq. 3 to be able to neutralize non-linearly encoded concepts while preserving the relative tractability and interpretability of the linear methods. A natural manner through which we can achieve these goals is by kernelization (Shawe-Taylor and Cristianini, 2004; Hofmann et al., 2008).
The underlying assumption motivating kernel methods is that the features needed for the task live in a reproducing kernel Hilbert space (RKHS; Aronszajn, 1950). At an intuitive level, an RKHS allows us to extend some of results from linear algebra to potentially infinite-dimensional vector spaces (Canu and Smola, 2006). The main technical contribution of this paper is the derivation of a kernelized version of the linear adversarial game presented in Eq. 3. We perform this derivation in this section after providing some background on kernel methods. We show that the resulting kernelized game is both non-convex and too computationally heavy to solve directly. Thus, we introduce a Nyström approximation that results in an efficient and light-weight formulation, given in Eq. 10, that is still able to isolate concepts encoded non-linearly.
3.1 Background: Kernel Methods
Kernel methods are based on reproducing kernel Hilbert spaces (RKHS). Without going into the technical details, an RKHS is a space of “nice” functions equipped with a kernel (Yosida, 2012). A kernel $\kappa(\cdot,\cdot)\geq 0$ is a similarity measure that generalizes a positive definite matrix. When we have a kernel over an RKHS, then the kernel corresponds to the dot product of a feature map, i.e., $\kappa(\mathbf{x},\mathbf{y})=\boldsymbol{\Phi}(\mathbf{x})^{\top}\boldsymbol{\Phi}(\mathbf{y})$. This insight gives us a natural manner to construct kernels. For instance, the linear kernel $\kappa(\mathbf{x},\mathbf{y})=\mathbf{x}^{\top}\mathbf{y}$ corresponds to the standard dot product in Euclidean space. The degree-2 polynomial kernel $\kappa(\mathbf{x},\mathbf{y})={(\gamma\mathbf{x}^{\top}\mathbf{y}+\alpha)}^{2}$ corresponds to a dot product in a six-dimensional feature space if $\mathbf{x},\mathbf{y}\in\mathbb{R}^{2}$.
Kernels are more general than positive definite matrices in that they can exist in infinite-dimensional spaces. For example, the Gaussian kernel $\exp\left(-\gamma||\mathbf{x}-\mathbf{y}||_{2}^{2}\right)$ is infinite-dimensional. [^3] One way to see this is to show a Gram matrix of any size $N$ will always have full rank. However, for any finite set of points $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$, we can construct a Gram matrix $K\in\mathbb{R}^{N\times N}$ where $K_{nm}=\kappa(\mathbf{x}_{n},\mathbf{x}_{m})$ encodes the similarity between $\mathbf{x}_{n}$ and $\mathbf{x}_{m}$. The matrix $K$ is guaranteed to be positive definite. Kernels are useful because they allow us to implicitly learn functions in a potentially infinite-dimensional RKHS without materializing that space.
3.2 A Kernelized Minimax Game
Inspection of the linear adversarial game in Eq. 1 reveals that both the adversary and the predictor interact with the input only via an inner product. Thus, the game can be kernelized by replacing the inner product with a kernel operation. We establish this kernelization by first proving the following representer theorem-like lemma, which shows that $\boldsymbol{w},{\boldsymbol{\theta}}$ can be written in terms of spans of the projected training set.
Lemma 1.
(Minimax Game Representer Theorem) Let $\mathcal{H}$ be a reproducing kernel Hilbert space with canonical feature map $\boldsymbol{\Phi}:\mathbb{R}^{D}\rightarrow\mathcal{H}$, i.e., $\boldsymbol{\Phi}(\mathbf{x})=\kappa(\mathbf{x},\cdot)$. Consider the game:
$\max_{\boldsymbol{w}\in\mathcal{H}}\min_{{\boldsymbol{\theta}}\in\mathcal{H}}\sum_{n=1}^{N}\ell\left(y_{n},\,\langle{\boldsymbol{\theta}},\mathrm{P}_{\boldsymbol{w}}^{\perp}\,\boldsymbol{\Phi}(\mathbf{x}_{n})\rangle\right)$ (5)
[…]
Proof.
See § A.1 for the proof. ∎
Having expressed $\boldsymbol{w},{\boldsymbol{\theta}}$ as a function of the training data, we can proceed to proving the kernelization of the adversarial game.
Lemma 2.
[…]
$-\frac{\boldsymbol{\alpha}^{\top}K^{(m)}(\mathbf{z})\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\bigg{)}$ (6)
where $K^{(m)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$.
Proof.
See § A.2 for the proof. ∎
Lemma 2 suggests the following form of the kernelized game given in Eq. 5:
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{N}}\,\max_{\boldsymbol{\alpha}\in\mathbb{R}^{N}}\,\sum_{n=1}^{N}\ell\Bigg{(}y_{n},\,$ $\sum_{m=1}^{N}\beta_{m}\bigg{(}\kappa(\mathbf{x}_{m},\mathbf{z}_{n})$
$-\frac{\boldsymbol{\alpha}^{\top}K^{(m,n)}\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\bigg{)}\Bigg{)}$ (7)
where we define $K^{(m,n)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z}_{n})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$. In contrast to Eq. 5, all computations in § 3.2 are in Euclidean space.
Theorem 1.
The reproducing kernel Hilbert space game Eq. 5 attains the same local optima as § 3.2.
Proof.
First, plug the result of Lemma 2 into Eq. 5. Optimality follows by Lemma 1, the representer lemma. ∎
Now, we turn to the runtime of the game.
Proposition 1.
The objective in § 3.2 can be computed in $\mathcal{O}(N^{4})$ time.
Proof.
Assuming that $\kappa(\cdot,\cdot)$ may be computed in $\mathcal{O}(1)$, computing $K^{(m,n)}$ takes $\mathcal{O}(N^{2})$ time. We have to do $\mathcal{O}(N^{2})$ such computations, which results in $\mathcal{O}(N^{4})$ time. Note that $K$ may be pre-computed once in $\mathcal{O}(N^{2})$ time. Thus, $K^{(m,n)}$ is the bottleneck, so the whole algorithm takes $\mathcal{O}(N^{4})$ time.∎
There are two problems with naïvely using the formulation in § 3.2. First, and similarly to Eq. 1, the problem is not convex–concave due to the optimization over $\boldsymbol{\alpha}$, which implicitly defines an orthogonal projection matrix of rank 1. Second, the evaluation time of § 3.2 is $\mathcal{O}\left(N^{4}\right)$, as argued in Proposition 1, which makes using a training set larger than a few hundred examples infeasible. We solve both of these computational issues with the Nyström approximation.
3.3 The Nyström Approximation
The general idea behind the Nyström method (Nyström, 1930) is to calculate a low-rank approximation of the kernel matrix. It is a commonly used technique for improving the runtime of kernel methods (Williams and Seeger, 2000).
3.3.1 Convexifying the Objective
Consider the Gram matrix $K\in\mathbb{R}^{N\times N}$ where $K_{nm}=\kappa(\mathbf{x}_{n},\mathbf{x}_{m})$, $\mathbf{x}_{n}$ is the $n^{\text{th}}$ training representation, and $\mathbf{x}_{m}$ is the $m^{\text{th}}$ training representation. We start with the eigendecomposition of $K=U\Sigma U^{\top}=U\sqrt{\Sigma}\sqrt{\Sigma}U^{\top}$, which can be computed in $\mathcal{O}\left(N^{3}\right)$ time (Golub and Van Loan, 2013). We are justified in taking the square root of $\Sigma$ because $K$ is necessarily positive definite. Now, we define an approximate feature map for observations $\mathbf{x}_{n}$ in the training data using the eigenvalues and vectors:
$\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\left(U\sqrt{\Sigma}\right)_{n}$ (8)
To compute the features for representations $\mathbf{x}$ not in the training data, we use the following:
$\widetilde{\boldsymbol{\Phi}}\left(\mathbf{x}\right)\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{n=1}^{N}\kappa(\mathbf{x},\mathbf{x}_{n})\widetilde{\boldsymbol{\Phi}}(\mathbf{x}_{n})$ (9)
which is an average, weighted by the kernel $\kappa(\cdot,\cdot)$, of the features obtained during training. Plugging Eq. 8 into Eq. 3 yields:
$\min_{{\boldsymbol{\theta}}\in\Theta}\,\max_{P\in\mathcal{F}_{k}}\,\sum_{n=1}^{N}\ell\Big{(}y_{n},\,\langle{\boldsymbol{\theta}},P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})\rangle\Big{)}$ (10)
which is identical to the linear game in Eq. 3, except that it uses the transformed features $\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})$ to approximate the true feature map $\boldsymbol{\Phi}$.
While neither the game in Eq. 3 nor the version of Eq. 3 with the Nyström approximation given in Eq. 10 is convex–concave, we are still able to approximately solve the game in practice. The runtime of a gradient-based optimization procedure for this game over $T$ epochs is now $\mathcal{O}\left(TN^{2}+N^{3}\right)$. [^5] Calculating the Nyström features entails computing the eigendecomposition of the Gram matrix, which is $\mathcal{O}\left(N^{3}\right)$. Then, we calculate the objective over all $N$ examples for $T$ epochs, which is $\mathcal{O}\left(TN^{2}\right)$. While this is an improvement over $\mathcal{O}\left(TN^{4}\right)$, it is still not fast enough to be of practical use.
3.3.2 Improving the Runtime
The runtime bottleneck of the game in Eq. 10 is the $\mathcal{O}\left(N^{3}\right)$ time it takes to compute the eigendecomposition of the Gram matrix $K$. Under the assumption that $\mathrm{rank}\left(K\right)=L$, we can improve this bound to $\mathcal{O}\left(L^{3}+L^{2}N\right)$ (Drineas and Mahoney, 2005). For the case that $L\ll N$, this is a substantial improvement. Moreover, it implies that the approximation feature map $\widetilde{\boldsymbol{\Phi}}(\mathbf{x}_{n})\in\mathbb{R}^{L}$. After $T$ steps of optimization, the runtime is now $\mathcal{O}\left(TL^{2}+L^{3}+L^{2}N\right)$, which is fast enough to be useful in practice. A natural question to ask is what happens if we apply the Nyström approximation, thereby assuming $\mathrm{rank}\left(K\right)=L$, but in practice $\mathrm{rank}\left(K\right)>L$? In this case, we are effectively computing a low-rank approximation of the kernel matrix. Several bounds on the accuracy of this approximation have been proven in the literature (Drineas and Mahoney, 2005; Jin et al., 2013; Nemtsov et al., 2016); we refer the reader to these works for more details on the approximation error.
3.4 Pre-image Mapping
After solving the game in Eq. 10, we obtain a projection matrix $P$ that neutralizes the concept inside an (approximated) RKHS. In other words, we have a function $r(\cdot)$ that prevents a classifier from predicting a concept label from the representation $\widetilde{\boldsymbol{\Phi}}(\mathbf{x}_{n})$. However, for many applications, we want a version of the input $\mathbf{x}_{n}$ in the original space with the concept neutralized, i.e., we want $r(\mathbf{x}_{n})$. Neutralization in the original space requires solving the pre-image problem (Mika et al., 1998). In the case of Nyström features, we seek a mapping $P\widetilde{\boldsymbol{\Phi}}(\mathbf{x})\mapsto\mathbf{x}$, which is a mapping from $\mathbb{R}^{L}$ to $\mathbb{R}^{D}$, that projects the neutralized features back into the input space.
In practice, this task can also be performed via a mapping $\mathbf{x}\mapsto\mathbf{x}$ from $\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$, which learns to reproduce in the input space the transformation that $P$ performs in the RKHS. We choose the latter approach, and train a multilayer perceptron (MLP) $f_{\lambda}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. [^6] This mapping may not always be a function because there may be more than one point in the RKHS that corresponds to a point in the original space. To estimate the parameters $\lambda$ of the MLP $f_{\lambda}(\cdot)$, we optimize a two-termed objective over all points $\mathbf{x}_{n}$:
$\operatorname*{\mathop{\text{argmin}}}_{\lambda\in\Lambda}$ ${\left|\left|P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})-\boldsymbol{\boldsymbol{\widetilde{\Phi}}}(f_{\lambda}(\mathbf{x}_{n}))\right|\right|_{2}^{2}}$ (11)
$\quad+\left|\left|(I-P)\boldsymbol{\widetilde{\Phi}}(f_{\lambda}(\mathbf{x}_{n}))\right|\right|_{2}^{2}$
where $\Lambda$ is the parameter space. The first term encourages $f_{\lambda}(\cdot)$ to perform the same transformation in the input space as $P$ does in the RKHS. The second term ensures that $P$ has no effect on the RKHS features computed on the neutralized $f_{\lambda}(\mathbf{x}_{n})$.
4 Experimental Setup
In § 3, we established an algorithm that allows us to attempt kernelized concept erasure of non-linearly encoded concepts. To summarize, this method requires first solving the game in Eq. 10 for a chosen neutralizing kernel, then training a pre-image network according to Eq. 11 to obtain neutralized representations in the input space.
We hypothesize that a non-linearly encoded concept can be exhaustively removed after mapping into the right RKHS. In order for this to hold, the neutralized representations must satisfy two conditions: Adversaries using non-linear classifiers should not be able to predict the erased concept from these representations, and these representations should preserve all other information encoded in them prior to erasure. With binary gender as our non-linearly encoded concept, we conduct several experiments testing both conditions. Before presenting our results, we lay out our experimental setup (see § A.3 for more details).
Data.
We run our main experiments on the identification and erasure of binary gender in static GloVe representations (Pennington et al., 2014). We focus on Ravfogel et al.’s (2020) dataset, where word representations are coupled with binary labels indicating whether they are male-biased or female-biased. As a preprocessing step, we normalize the GloVe representations to have unit norm. For an extrinsic evaluation of our method on a main task (profession prediction), we use the Bias-in-Bios dataset of De-Arteaga et al. (2019), which consists of a large set of short biographies annotated for both gender and race. Following Ravfogel et al. (2022), we embed each biography using the [CLS] representation of pre-trained BERT.
Kernels.
We consider the following kernels:
  • Poly : $\kappa(\mathbf{x},\mathbf{y})={(\gamma\mathbf{x}^{\top}\mathbf{y}+\alpha)}^{\texttt{d}}$
  • RBF : $\kappa(\mathbf{x},\mathbf{y})=\exp\left(-\gamma||\mathbf{x}-\mathbf{y}||_{2}^{2}\right)$
  • Laplace : $\kappa(\mathbf{x},\mathbf{y})=\exp\left(-\gamma||\mathbf{x}-\mathbf{y}||_{1}\right)$
  • Linear : $\kappa(\mathbf{x},\mathbf{y})=\mathbf{x}^{\top}\mathbf{y}$
  • Sigmoid : $\kappa(\mathbf{x},\mathbf{y})=\mathrm{tanh}(\gamma\mathbf{x}^{\top}\mathbf{y}+\alpha)$
  • Multiple : a convex combination of the above kernels. We consider the following two methods for combining kernels:
    • EasyMKL : a convex combination learned with the EasyMKL algorithm (Aiolli and Donini, 2015) targeted for gender prediction. [^7] We use the implementation of MKLpy with the default parameters.
    • UniformMK : a uniform combination of all kernels.
We experiment with different values for the hyperparameters $\gamma>0$, $\alpha>0$ and $\texttt{d}>0$ (see § A.3 for details). We use $L=1024$-dimensional vectors for the Nyström approximation.
Reported metrics.
Each result is reported as the mean $\pm$ standard deviation, computed across four runs of the experiment with random restarts.
Solving the minimax game.
We solve the relaxed adversarial game given in Eq. 10 by alternate gradient-based optimization (Goodfellow et al., 2016). Concretely, we alternate between updating the predictor’s parameters ${\boldsymbol{\theta}}$ and the projection matrix $P$. Updates to ${\boldsymbol{\theta}}$ are performed with gradient descent, and updates to $P$ are performed with gradient ascent, including a projection onto the Fantope to ensure that the constraint is met. For the Fantope projection step, we use Vu et al.’s (2013) algorithm, the details of which are restated in Ravfogel et al. (2022).
Pre-image calculation.
As our pre-image network $f_{\lambda}(\cdot)$, we use an MLP with two hidden layers of sizes 512 and 300, respectively. We use layer normalization after each hidden layer and ReLU activations. See § A.3.1 for basic empirical validation of the pre-image calculation procedure.
5 Effect on Concept Encoding
In this section, we pose the exhaustive RKHS hypothesis: The hypothesis that binary gender can be exhaustively removed when the representations are mapped into the right RKHS. That is, there exists a unique kernel such that, for any choice of non-linear predictor, the adversary cannot recover gender information from the pre-image representations obtained via our special kernel. As a baseline, we note that gender prediction accuracy on the original representations, prior to any intervention, is above 99% with every kernel, including the linear kernel. This means that the gender concept is linearly separable in the original input space. In this context, we conduct the following experiments on gender neutralization.
Same adversary.
We start by calculating the neutralized pre-images for each kernel type, and then apply the same kernel adversary to recover gender information. This experiment tests whether we can protect against the same kernel adversary.
Transfer between kernels.
To directly test the exhaustive RKHS hypothesis, we calculate the neutralized pre-image representations with respect to a neutralizing kernel, and then use a different adversarial kernel to recover gender. For instance, we calculate neutralized pre-image representations with respect to a polynomial kernel, and then predict gender using a Laplace kernel, or a polynomial kernel with different hyperparameters.
Biased associations.
Gender can manifest itself in the pre-image representations via biased associations, even when gender is neutralized according to our adversarial test. To assess the impact of our intervention on this notion of gender encoding, we run the WEAT test (Islam et al., 2017) on the neutralized pre-image representations.
Type Accuracy
Poly 0.59 $\pm$ 0.15
RBF 0.69 $\pm$ 0.16
Laplace 0.75 $\pm$ 0.11
Linear 0.54 $\pm$ 0.02
Sigmoid 0.49 $\pm$ 0.00
EasyMKL 0.69 $\pm$ 0.01
UniformMK 0.49 $\pm$ 0.00
Table 1: Gender prediction accuracy from the neutralized pre-image representations when using the same kernel for neutralization and recovery. Numbers are averages over hyperparameters of each kernel and over four randomized runs of the experiment.
Poly RBF Laplace Linear Sigmoid EasyMKL UniformMK MLP
Poly 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.93 $\pm$ 0.00 0.55 $\pm$ 0.01 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.49 $\pm$ 0.00 0.97 $\pm$ 0.00
RBF 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.93 $\pm$ 0.00 0.59 $\pm$ 0.02 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.49 $\pm$ 0.00 0.97 $\pm$ 0.00
Laplace 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.61 $\pm$ 0.01 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.59 $\pm$ 0.02 0.97 $\pm$ 0.00
Linear 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.93 $\pm$ 0.00 0.54 $\pm$ 0.02 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.49 $\pm$ 0.00 0.97 $\pm$ 0.00
Sigmoid 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.89 $\pm$ 0.01 0.65 $\pm$ 0.03 0.49 $\pm$ 0.00 0.97 $\pm$ 0.00 0.64 $\pm$ 0.03 0.97 $\pm$ 0.00
EasyMKL 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.57 $\pm$ 0.03 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.49 $\pm$ 0.00 0.97 $\pm$ 0.00
UniformMK 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.01 0.58 $\pm$ 0.08 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.49 $\pm$ 0.00 0.97 $\pm$ 0.00
Table 2: Evaluation of the neutralized pre-image representations when using a different kernel for neutralization and recovery. The neutralizing kernel is presented in the rows, and the kernel adversary used for recovery in the columns. Note that on the diagonal, the hyperparameters of the same kernel family differ between the neutralizing kernel and the adversary that recovers gender form the pre-image representations. See § A.7 for a breakdown by neutralizing kernel hyperparameters and for details on the hyperparameters of the kernel adversaries.
Kernel WEAT’s $d$ WEAT’s $p$-value
Poly 0.74 $\pm$ 0.01 0.08 $\pm$ 0.00
RBF 0.74 $\pm$ 0.00 0.08 $\pm$ 0.00
Laplace 0.71 $\pm$ 0.03 0.09 $\pm$ 0.01
Linear 0.74 $\pm$ 0.00 0.08 $\pm$ 0.00
Sigmoid 0.75 $\pm$ 0.02 0.08 $\pm$ 0.01
EasyMKL 0.73 $\pm$ 0.00 0.08 $\pm$ 0.00
UniformMK 0.73 $\pm$ 0.00 0.08 $\pm$ 0.00
Original 1.56 0.000
Table 3: WEAT results on pre-image representations. Numbers are averages over hyperparameters of each kernel.
5.1 Pre-image Gender Recovery: Same Adversary
In Table 1, we report average gender prediction accuracy on the neutralized pre-images for the case where the neutralizing and adversarial kernels are of the same type and share the same hyperparameters. [^8] We use the Nyström approximation for neutralization, but predict gender from the pre-image representations using SVM classifiers that use the true kernel. Numbers are averages over the results of all hyperparameter values of each kernel. See § A.4 for the full results. As can be seen, for most—but not all—kernels, we effectively hinder the ability of the non-linear classifier to predict gender.
5.2 Pre-image Gender Recovery: Transfer Between Kernels
In Table 2, we report average gender prediction accuracy on the neutralized pre-images for the case where the neutralizing and adversarial kernels are different. Numbers are averages over several different hyperparameter settings for the neutralizing kernel, while the adversarial kernel hyperparameters are fixed as detailed in § A.7. Also, see the appendix for a full breakdown by neutralizing kernel hyperparameters. Table 2 allows us to test the exhaustive RKHS hypothesis. Under this hypothesis, we would expect to see that for at least one neutralizing kernel, no non-linear classifers are able to accurately recover gender. For a thorough test of the hypothesis, we introduce an MLP with a single hidden layer as an additional adversary.
Remarkably, we observe a complete lack of generalization of our concept erasure intervention to other types of non-linear predictors. In particular, no neutralizing kernel significantly hinders the ability of an MLP with a single hidden layer to predict gender: the MLP always recovers the gender labels with an accuracy of 97%. Furthermore, concept erasure does not transfer between different kernel types, and even between kernels of the same family with different hyperparameter settings. For instance, when using a polynomial neutralizing kernel, we protect against a polynomial adversarial kernel with the same parameters (mean accuracy of 59% in Table 1), but not against a polynomial adversarial kernel with different hyperparameters (98% mean accuracy for Poly in Table 2).
We do see transfer to sigmoid, and—to a lesser degree—linear kernel adversaries, with a classification accuracy of 54–65% for the linear adversary, and 49% for the sigmoid adversary. Surprisingly, the sigmoid kernel seems weaker than the linear kernel, achieving a near-random accuracy of 49%. The results do not show evidence of a proper hierarchy in the expressiveness of the different kernels, and convex combinations of the different kernels—either learned (EasyMKL) or uniform (UniformMK)—do not provide better protection than individual kernels.
In short, while we are able to effectively protect against the same kernel, transfer to different kernels is non-existent. This result does not support the exhaustive RKHS hypothesis: We do not find a single RKHS that exhaustively encodes the binary gender concept.
5.3 Effect on Gendered Word Associations
In the case where the concept of gender cannot be recovered by an adversary, binary gender could still manifest itself in more subtle ways. For instance, it may be the case that the names of gender-biased professions, such as STEM fields, are closer in representation space to male-associated words than to female-associated words. We aim to measure the extent to which our neutralized pre-image representations exhibit this measure of gender bias.
Evaluation.
To quantify bias associations, Islam et al. (2017) propose the WEAT word association test. This test measures WEAT’s d, a statistic that quantifies the difference in similarity between two sets of gendered words (e.g., male first names and female first names) and two sets of potentially biased words (e.g., stereotypically male and stereotypically female professions). [^9] It asks, for example, whether the degree to which male first names are closer to scientific terms versus artistic terms is significantly different from the same quantity calculated for female first names. We repeat the experiments of Gonen and Goldberg (2019) and Ravfogel et al. (2020). Following Gonen and Goldberg (2019), we represent the male and female groups with names commonly associated with males and females, rather than with explicitly gendered words (e.g., pronouns). Three tests evaluate the association between name groups and i) career and family-related words; ii) art and mathematics-related words; and iii) names of artistic and scientific fields. Successful neutralization of gender would imply that these word groups are less closely associated in the pre-image representations.
Results.
In Table 3, we report the test statistic and the $p$ value for the third test using the names of scientific and artistic fields to represent gender-biased words. [^10] See § A.6 for results of the other two biased word association tests. For all kernels, we observe a significant drop in the test statistic, from the original value of 1.56 to around 0.74. This suggests that the intervention significantly decreases the association between female and male names and stereotypically biased words. Notably, the reduction is similar for all kernels, including the linear one. While non-linear erasure is more effective in neutralizing gender against adversarial recovery, linear and non-linear methods perform equally well according to this bias association test. This finding highlights the importance of measuring different manifestations of a concept when using concept neutralization as a bias mitigation method.
6 Negative Impact on the Representations
Our method has shown a satisfactory ability to prevent the same kernel from recovering the concept. However, does erasure remove too much information? [^11] It is trivial to neutralize a concept by completely zeroing out the representations. As previously stated, our intervention should erase a concept without altering any of the other information encoded in the original representations. In this section, we evaluate whether the non-gender related semantic content of the original representations is preserved in our neutralized pre-images. We do so via the following tests: i) an intrinsic evaluation of the semantic content of the neutralized pre-image word representation space, and ii) an extrinsic evaluation of our method when applied to contextualized word representations for a profession prediction task, measuring the extent to which we hinder a model’s ability to perform the main task.
6.1 Intrinsic Evaluation of Damage to Semantic Content
To measure the influence of our method on the semantics encoded in the representation space, we use SimLex-999 (Hill et al., 2015), an annotated dataset of word pairs with human similarity scores for each pair. First, we calculate the cosine similarity between the representations of each pair of words using the original representations. Then, we repeat this calculation for each type of kernel using the pre-image representations. Finally, we measure the correlation between the similarity of words in representation space and human similarity scores, before and after intervention. The original correlation is $0.400$, and it is left nearly unchanged by any of the kernel interventions, yielding values between $0.387$ and $0.396$. To qualitatively demonstrate the absence of negative impact, we show in § A.5 that the nearest neighbors of randomly sampled words do not change significantly after gender erasure.
6.2 Extrinsic Evaluation on Contextualized Representations
The previous experiments focused on the influence of concept neutralization on uncontextualized representations. Here, we apply our concept neutralization method on contextualized BERT representations and assess its effect on profession prediction. We embed each biography in the dataset of De-Arteaga et al. (2019) using the [CLS] representation of pre-trained BERT, and apply our method using only the RBF kernel. After collecting the pre-image representations, we train a linear classifier on the main task of profession prediction.
Results.
Averaged over different hyperparameter settings of the RBF kernel, we achieve a profession prediction accuracy after neutralization of $74.19\pm 0.056\%$. For reference, prediction accuracy using the original BERT representations is $76.93\%$. This suggests that the pre-images still encode most of the profession information, which is largely orthogonal to the neutralized gender information.
7 Discussion
We have demonstrated that in the case where the neutralizing kernel and the adversarial kernel are the same, we are able to neutralize a non-linearly encoded concept reasonably well. We have also shown that our method neutralizes gender in a comprehensive manner, without damaging the representation. However, this neutralization does not transfer to different non-linear adversaries, which are still able to recover gender.
While the lack of transfer to other non-linear predictors may seem surprising, one should keep in mind that changing the kernel type, or changing kernel hyperparameters, results in a different implicit feature mapping. Even the features defined by a linear kernel are not a proper subset of the features defined by a polynomial kernel of degree 2.121212Features defined by a linear kernel can be a proper subset of the features defined by a polynomial kernel if there is a fixed coordinate in all examples prior to the kernel mapping. As such, removing the features which make the concept of interest linearly separable in one RKHS does not necessarily prevent a classifier parameterized by another kernel or an MLP from predicting the concept. In the context of gender erasure, these results suggest that protection against a diverse set of non-linear adversaries remains an open problem.
8 Conclusion
We propose a novel method for the identification and erasure of non-linearly encoded concepts in neural representations. We first map the representations to an RKHS, before identifying and neutralizing the concept in that space. We use our method to empirically assess the exhaustive RKHS hypothesis: We hypothesize that there exists a unique kernel that exhaustively identifies the concept of interest. We find that while we are able to protect against a kernel adversary of the same type, this protection does not transfer to different non-linear classifiers, thereby contradicting to the RKHS hypothesis. Exhaustive concept erasure and protection against a diverse set of non-linear adversaries remains an open problem.
Limitations
The empirical experiments in this work involve the removal of binary gender information from pre-trained representations. We note the fact that gender is a non-binary concept as a major limitation of our work. This task may have real-world applications, in particular relating to fairness. We would encourage readers to be careful when attempting to deploy methods such as the one discussed in this paper. Regardless of any proofs, one should carefully measure the effectiveness of the approach in the context in which it is to be deployed. Please consider, among other things, the exact data to be used, the fairness metrics under consideration, and the overall application.
We urge practitioners not to regard this method as a solution to the problem of bias in neural models, but rather as a preliminary research effort toward mitigating certain aspects of the problem. Unavoidably, the datasets we use do not reflect all the subtle and implicit ways in which gender bias is manifested. As such, it is likely that different forms of bias still exist in the representations following the application of our method.
Ethical Concerns
We do not foresee ethical concerns with this work.
Acknowledgements
The authors sincerely thank Clément Guerner for his thoughtful and comprehensive comments and revisions to the final version of this work. This project received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, grant agreement No. 802774 (iEXTRACT). Ryan Cotterell acknowledges Google for support from the Research Scholar Program.
Appendix A Related Work
Mitigation of Gender Bias.
The identification of linear subspaces that encode binary gender has attracted considerable research interest (Bolukbasi et al., 2016; Gonen and Goldberg, 2019; Dev and Phillips, 2019; Ravfogel et al., 2020). While bias mitigation is a central use case of concept erasure, concept subspaces have been applied to a number of tasks. Concept subspaces have been used to analyze the content of neural representations, e.g., for causal analysis (Elazar et al., 2021; Ravfogel et al., 2021), for analyzing the geometry of the representation space (Celikkanat et al., 2020; Gonen et al., 2020; Hernandez and Andreas, 2021), and for concept-based interpretability (Kim et al., 2018).
Kernelization of Linear Methods.
The kernelization of linear machine learning algorithms is a common practice, and has many use cases, such as the kernelized perceptron (Aizerman et al., 1964) and kernel PCA (Schölkopf et al., 1997). White et al. (2021) proposed a kernelization of a structural probe that extracts syntactic structure from neural representations. Vargas and Cotterell (2020) proposed a kernelization of the PCA-based bias mitigation method of Bolukbasi et al. (2016), and found that it does not improve on the linear mitigation procedure. Since the effectiveness of this method has been questioned (Gonen and Goldberg, 2019), we consider a more principled and well-motivated approach for the identification and neutralization of the concept subspace. Sadeghi et al. (2019) proposed a kernelization of an alternative, regression-based linear adversarial objective, which is not limited to orthogonal projections. Our formulation is different in that it considers any linear model, and is restricted to the neutralization of linear subspaces via projection. This makes our method potentially less expressive, but more interpretable.
A.1 A Representer Lemma for Kernelized Minimax Games
See 1
Proof.
For brevity, first notice we can re-express the objective as
$\max_{\boldsymbol{w},\in\mathcal{H}}\min_{{\boldsymbol{\theta}}\in\mathcal{H}}\sum_{n=1}^{N}\ell\left(y_{n},\,\langle{\boldsymbol{\theta}},\left(\mathrm{I}-\frac{{\boldsymbol{w}\boldsymbol{w}^{\top}}}{\boldsymbol{w}^{\top}\boldsymbol{w}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})\rangle\right)$ (12)
We will show that both $\boldsymbol{w}$ and ${\boldsymbol{\theta}}$ can be expressed as a linear combination of terms from the training data without losing expressive power. Now, decompose $\boldsymbol{w}$ as follows: $\boldsymbol{w}=\boldsymbol{w}_{U}+\boldsymbol{w}_{\perp U}$, where we represent $\boldsymbol{w}$ as the sum of $\boldsymbol{w}$ projected onto $U$ and onto its orthogonal complement $U_{\perp}$. Now, note that for any element of the training data $\boldsymbol{\Phi}(\mathbf{x}_{n})$, we have
$\left(\mathrm{I}-\frac{\boldsymbol{w}\boldsymbol{w}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})$ $=\left(\mathrm{I}-\frac{(\boldsymbol{w}_{U}+\boldsymbol{w}_{\perp U})(\boldsymbol{w}_{U}+\boldsymbol{w}_{\perp U})^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})$ (13a)
$=\left(\mathrm{I}-\frac{\boldsymbol{w}_{U}\boldsymbol{w}_{U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}-\underbrace{\frac{\boldsymbol{w}_{U}\boldsymbol{w}_{\perp U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}}_{=0}-\underbrace{\frac{\boldsymbol{w}_{\perp U}\boldsymbol{w}_{U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}}_{=0}-\frac{\boldsymbol{w}_{\perp U}\boldsymbol{w}_{\perp U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})$ (13b)
$=\left(\mathrm{I}-\frac{\boldsymbol{w}_{U}\boldsymbol{w}_{U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}-\frac{\boldsymbol{w}_{\perp U}\boldsymbol{w}_{\perp U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})$ (13c)
$=\boldsymbol{\Phi}(\mathbf{x}_{n})-\frac{\boldsymbol{w}_{U}\boldsymbol{w}_{U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\boldsymbol{\Phi}(\mathbf{x}_{n})-\frac{\boldsymbol{w}_{\perp U}\boldsymbol{w}_{\perp U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (13d)
$=\boldsymbol{\Phi}(\mathbf{x}_{n})-\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}$ (13e)
However, we have that $\boldsymbol{\Phi}(\mathbf{x}_{n})-\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}$ is in $U$. Likewise, we can decompose ${\boldsymbol{\theta}}={\boldsymbol{\theta}}_{U}+{\boldsymbol{\theta}}_{\perp U}$. Further manipulation reveals
$\Big{\langle}{\boldsymbol{\theta}},\left(\mathrm{I}-{\boldsymbol{w}\boldsymbol{w}^{\top}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})\Big{\rangle}={\boldsymbol{\theta}}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})-{\boldsymbol{\theta}}_{U}^{\top}\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}$ (14)
Thus, for any ${\boldsymbol{\theta}},\boldsymbol{w},\in\mathcal{H}$ there exists a ${\boldsymbol{\theta}}_{U},\boldsymbol{w}_{U}\in U$ that yields the same value of the objective as ${\boldsymbol{\theta}},\boldsymbol{w}$. Now, we can parameterize ${\boldsymbol{\theta}}_{U}$ and $\boldsymbol{w}_{U}$ as
$\boldsymbol{w}_{U}$ $=\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (15)
${\boldsymbol{\theta}}_{U}$ $=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (16)
for real coefficients $\boldsymbol{\alpha}\in\mathbb{R}^{N}$ and $\boldsymbol{\beta}\in\mathbb{R}^{N}$. We conclude that for any local optimum ${\boldsymbol{\theta}}^{*},\boldsymbol{w}^{*}\in\mathcal{H}$, the projection of ${\boldsymbol{\theta}}^{*}$ and $\boldsymbol{w}^{*}$ onto $U$ yields a local optimum with the same value. ∎
We note that under regularity conditions, i.e., certain compactness and convexity restrictions over the feasible sets and the loss function, the min and the max can be swapped as per the celebrated Von Neumann–Fan minimax theorem (Fan, 1953). For the aforementioned reasons, we believe Lemma 1 justifies the parameterizations used for ${\boldsymbol{\theta}},\boldsymbol{w}$, e.g., Eq. 19.
A.2 Kernelization of the Minimax Game
We show that the game Eq. 1 can be kernelized for the case $k=1$, i.e., a setting where the matrix $\mathcal{P}_{k}$ removes a one-dimensional subspace. Specifically, we will show that the product $\langle{\boldsymbol{\theta}},\mathrm{P}\boldsymbol{\Phi}(\mathbf{x}_{n})\rangle$ in Eq. 1 can be expressed as a function of the kernel $\kappa(\cdot,\cdot)$.
See 2
Proof.
The projection onto the orthogonal complement of $\boldsymbol{w}=\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ is defined as the following
$\mathrm{P}_{\boldsymbol{w}}^{\perp}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{I}-\frac{\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{i})}\Big{)}\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}}\Big{)}}{\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}\Big{)}\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})\Big{)}}$ (17)
where $\mathrm{I}$ is the identity operator. Algebraic manipulation reveals
$\mathrm{P}_{\boldsymbol{w}}^{\perp}\boldsymbol{\Phi}(\mathbf{z})$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{I}\boldsymbol{\Phi}(\mathbf{z})-\frac{\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})}\Big{)}\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n^{\prime}})^{\top}}\Big{)}}{\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}\Big{)}\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})\Big{)}}\,\boldsymbol{\Phi}(\mathbf{z})$ (18a)
$=\boldsymbol{\Phi}(\mathbf{z})-\frac{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\mathbf{x}_{n})\boldsymbol{\Phi}(\mathbf{x}_{m})^{\top}}\boldsymbol{\Phi}(\mathbf{z})}{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}\boldsymbol{\Phi}(\mathbf{x}_{m})}}$ (18b)
$=\boldsymbol{\Phi}(\mathbf{z})-\frac{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\mathbf{x}_{n})\kappa(\mathbf{x}_{m},\mathbf{z})}}{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\kappa(\mathbf{x}_{n},\mathbf{x}_{m})}}$ (18c)
$=\boldsymbol{\Phi}(\mathbf{z})-\underbrace{\left(\frac{\sum_{m=1}^{N}{\alpha_{m}\kappa(\mathbf{x}_{m},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)}_{\in\mathbb{R}}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (18d)
$=\boldsymbol{\Phi}(\mathbf{z})-\left(\frac{\sum_{m=1}^{N}{\alpha_{m}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}$ (18e)
=$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\boldsymbol{\Phi}_{\text{proj}}(\mathbf{z})}$ (18f)
Now, consider an element of the reproducing kernel Hilbert space
${\boldsymbol{\theta}}=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (19)
Further algebraic manipulation reveals
$\Big{\langle}{\boldsymbol{\theta}},$ ${\boldsymbol{\Phi}_{\text{proj}}(\mathbf{z})}\Big{\rangle}=\Big{\langle}{\boldsymbol{\theta}},\boldsymbol{\Phi}(\mathbf{z})-\left(\frac{\sum_{n=1}^{N}{\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}\Big{\rangle}$ (20a)
$=\Big{\langle}\sum_{m=1}^{N}\beta_{m}\boldsymbol{\Phi}(\mathbf{x}_{m}),\boldsymbol{\Phi}(\mathbf{z})-\left(\frac{\sum_{n=1}^{N}{\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}\Big{\rangle}$ (20b)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\left(\frac{\sum_{n=1}^{N}{\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\kappa(\mathbf{x}_{m},\boldsymbol{w})$ (20c)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\frac{1}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\sum_{n=1}^{N}\sum_{m=1}^{N}\alpha_{n}\beta_{m}\kappa(\mathbf{x}_{n},\mathbf{z})\kappa(\mathbf{x}_{m},\boldsymbol{w})$ (20d)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})\kappa(\mathbf{x}_{m},\boldsymbol{w})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20e)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})\boldsymbol{\Phi}(\mathbf{x}_{m})^{\top}\left(\sum_{n^{\prime}=1}^{N}\alpha_{n^{\prime}}\boldsymbol{\Phi}(\mathbf{x}_{n^{\prime}})\right)\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20f)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\alpha_{n^{\prime}}\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})\boldsymbol{\Phi}(\mathbf{x}_{m})^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n^{\prime}})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20g)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\alpha_{n}\alpha_{n^{\prime}}\kappa(\mathbf{x}_{n},\mathbf{z})\kappa(\mathbf{x}_{m},\mathbf{x}_{n^{\prime}})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20h)
$=\sum_{m=1}^{N}\beta_{m}\left(\kappa(\mathbf{x}_{m},\mathbf{z})-\frac{\boldsymbol{\alpha}^{\top}K^{(m)}(\mathbf{z})\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)$ (20i)
where we define the following matrix component-wise: $K^{(m)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$. Eq. 20i can be evaluated without explicitly applying the kernel transformation $\boldsymbol{\Phi}$. In terms of notation, when we have $\Big{\langle}{\boldsymbol{\theta}},{\boldsymbol{\Phi}_{\text{proj}}}(\mathbf{z}_{n})\Big{\rangle}$, we write
$\sum_{m=1}^{N}\beta_{m}\left(\kappa(\mathbf{x}_{m},\mathbf{z}_{n})-\frac{\boldsymbol{\alpha}^{\top}K^{(m,n)}\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)$ (21)
where we define $K^{(m,n)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z}_{n})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$. This proves the result. ∎
A.3 Experimental setting
Data.
We conduct experiments on the uncased version of the GloVe representations, which are 300-dimensional, licensed under Apache License, Version 2.0. Following Ravfogel et al. (2020), to approximate the gender labels for the vocabulary, we project all representations on the $\overrightarrow{\text{he}}-\overrightarrow{\text{she}}$ direction, and take the $7,500$ most male-biased and female-biased words. Note that unlike (Bolukbasi et al., 2016), we use the $\overrightarrow{\text{he}}-\overrightarrow{\text{she}}$ direction only to induce approximate gender labels, but then proceed to measure the bias in various ways that go beyond neutralizing just the $\overrightarrow{\text{he}}-\overrightarrow{\text{she}}$ direction. We use the same train–dev–test split of Ravfogel et al. (2020), but discard the gender-neutral words (i.e., we cast the problem as a binary classification). We obtain training, evaluation, and test sets of sizes 7,350, 3,150 and 4,500, respectively. We perform four independent runs of the entire method for all kernel types, with different random seeds.
The kernelized minimax game.
For each kernel, we experiment with the following combinations of hyperparameter values:
  • Poly : $\texttt{d}\ \in\{2,3\}$; $\gamma\in\{0.05,0.1,0.15\}$; $\alpha\in\{0.8,1,1.2\}$.
  • RBF : $\gamma\in\{0.1,0.15,0.2\}$
  • Laplace : $\gamma\in\{0.1,0.15,0.2\}$
  • Sigmoid : $\alpha\in\{0,0.01\}$; $\gamma\in\{0.005,0.003\}$.
We approximate the kernel space using $L=1024$ Nyström landmarks. We run the adversarial game Eq. 10 for each of the kernel mappings we consider, by performing alternate minimization and maximization over ${\boldsymbol{\theta}}$ and $P$, respectively. As our optimization procedure, we use stochastic gradient descent with a learning rate of 0.08 and minibatches of size 256. We run for 35,000 batches, and choose the projection matrix $P$ which leads to the biggest decrease in the linear classification accuracy on the evaluation set. In all cases, we identify a matrix which decreases classification accuracy to near-random accuracy. All training is done on a single NVIDIA GeForce GTX 1080 Ti GPU.
Pre-image mapping.
We train an MLP with 2 hidden layers of sizes 512 and 300 to map the original inputs $\mathbf{x}_{n}$ to inputs which, after being mapped to kernel space, are close to the neutralized features. We use dropout of 0.1, ReLU activation and layer normalization after each hidden layer. We use a skip connection between the input and the output layer, i.e., we set the final output of the MLP to be the sum of its inputs and outputs. [^13] This is useful because we assume the gender information is not very salient in the original representations, since these encode much more information. With the skip connection, the MLP only needs to learn what to remove, rather than learning to generate plausible representations. We train for 15,000 batches of size 128 and choose the model that yields the lowest loss on the evaluation set.
Non-linear gender prediction.
We consider the following non-linear predictors: SVMs with different kernels, as well as an MLP with 128 hidden units and ReLU activations. We use the sklearn implementation of predictors. They are trained on the reconstructed pre-image of the training set, and tested on the reconstructed pre-image of the test set. Note that while in training we used an approximation of the kernel function, we predict gender from the pre-images using SVM classifiers that rely on the actual, exact kernel.
A.3.1 Pipeline Evaluation
In this appendix, we include sanity check experiments that aim to assess whether the minimax game Eq. 10 effectively removes linearly-present concepts from the non-linear kernel features, and whether the training of the pre-image network succeeds.
Concept erasure in kernel space.
Do we effectively neutralize the concept in the approximate kernel space? For each kernel, we solve the game Eq. 10 and use the final projection matrix $P$ to create neutralized features. We neutralize the features in RKHS by mapping $\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})\mapsto P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})$, and train a linear classifier to recover the gender labels from the neutralized representations. We get a classification accuracy of 50.59$\pm$0.04, very close to majority accuracy of 50.58. This suggests that the process is effective in protecting against the kernel which was applied in training (training a linear classifier on the kernel-transformed representations is equivalent to training a kernel classifier on the original representations). Notice, however, that we cannot test other non-linear kernel classifiers on these representations in a similar way: If the approximate kernel mapping $\widetilde{\boldsymbol{\Phi}}(\cdot)$ corresponds, for example, to a polynomial kernel, we cannot measure the success of an RBF kernel in recovering the bias information after the intervention without performing the pre-image mapping.
Pre-image mapping.
Our neutralization algorithm relies on calculating the pre-image of the kernel features after the intervention. To evaluate the quality of the pre-image mapping, we measure the relative reconstruction error $\left|\left|\frac{{P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})-\boldsymbol{\widetilde{\Phi}}(f(\mathbf{x}_{n}))}}{P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})}\right|\right|^{2}_{2}$ over all points $\mathbf{x}_{n}$ on the evaluation set. When averaged over all 4 seeds and the different kernels we experimented with, we get a reconstruction error of 1.81$\pm$1.61(range 0.45-7.94).
A.4 Gender Prediction from the Pre-image
In Table 4 we report the full evaluation results on the pre-image neutralized representations, where we have used the same kernel (and the same hyperparameters) for neutralization and gender recovery from the pre-image.
Kernel $\gamma$ $\alpha$ d WEAT’s d WEAT’s p-value Gender Acc.
Poly 0.05 0.8 2 0.73 $\pm$ 0.01 0.084 $\pm$ 0.002 0.49 $\pm$ 0.00
Poly 0.05 1 2 0.74 $\pm$ 0.01 0.080 $\pm$ 0.002 0.49 $\pm$ 0.00
Poly 0.05 1.2 2 0.74 $\pm$ 0.01 0.081 $\pm$ 0.002 0.49 $\pm$ 0.00
Poly 0.05 0.8 3 0.75 $\pm$ 0.01 0.077 $\pm$ 0.003 0.49 $\pm$ 0.00
Poly 0.05 1 3 0.74 $\pm$ 0.01 0.080 $\pm$ 0.003 0.49 $\pm$ 0.00
Poly 0.05 1.2 3 0.74 $\pm$ 0.01 0.081 $\pm$ 0.003 0.49 $\pm$ 0.00
Poly 0.1 0.8 2 0.74 $\pm$ 0.01 0.080 $\pm$ 0.003 0.49 $\pm$ 0.00
Poly 0.1 1 2 0.74 $\pm$ 0.00 0.080 $\pm$ 0.001 0.49 $\pm$ 0.00
Poly 0.1 1.2 2 0.74 $\pm$ 0.01 0.081 $\pm$ 0.002 0.49 $\pm$ 0.00
Poly 0.1 0.8 3 0.74 $\pm$ 0.00 0.081 $\pm$ 0.002 0.53 $\pm$ 0.01
Poly 0.1 1 3 0.73 $\pm$ 0.01 0.082 $\pm$ 0.004 0.65 $\pm$ 0.03
Poly 0.1 1.2 3 0.73 $\pm$ 0.01 0.084 $\pm$ 0.004 0.72 $\pm$ 0.01
Poly 0.15 0.8 2 0.73 $\pm$ 0.01 0.082 $\pm$ 0.003 0.56 $\pm$ 0.03
Poly 0.15 1 2 0.74 $\pm$ 0.02 0.081 $\pm$ 0.006 0.54 $\pm$ 0.02
Poly 0.15 1.2 2 0.73 $\pm$ 0.01 0.084 $\pm$ 0.004 0.55 $\pm$ 0.02
Poly 0.15 0.8 3 0.73 $\pm$ 0.01 0.082 $\pm$ 0.003 0.88 $\pm$ 0.02
Poly 0.15 1 3 0.73 $\pm$ 0.01 0.082 $\pm$ 0.003 0.92 $\pm$ 0.00
Poly 0.15 1.2 3 0.74 $\pm$ 0.01 0.079 $\pm$ 0.003 0.93 $\pm$ 0.00
RBF 0.1 - - 0.75 $\pm$ 0.01 0.078 $\pm$ 0.003 0.49 $\pm$ 0.01
RBF 0.15 - - 0.74 $\pm$ 0.01 0.079 $\pm$ 0.003 0.68 $\pm$ 0.03
RBF 0.2 - - 0.74 $\pm$ 0.01 0.081 $\pm$ 0.003 0.89 $\pm$ 0.01
Laplace 0.1 - - 0.72 $\pm$ 0.03 0.086 $\pm$ 0.008 0.62 $\pm$ 0.04
Laplace 0.15 - - 0.74 $\pm$ 0.05 0.080 $\pm$ 0.015 0.77 $\pm$ 0.05
Laplace 0.2 - - 0.67 $\pm$ 0.05 0.107 $\pm$ 0.020 0.88 $\pm$ 0.04
Linear - - - 0.74 $\pm$ 0.01 0.079 $\pm$ 0.004 0.54 $\pm$ 0.02
Sigmoid 0.005 0 - 0.78 $\pm$ 0.05 0.069 $\pm$ 0.014 0.49 $\pm$ 0.00
Sigmoid 0.005 0.01 - 0.73 $\pm$ 0.03 0.082 $\pm$ 0.010 0.49 $\pm$ 0.00
Sigmoid 0.003 0 - 0.73 $\pm$ 0.05 0.083 $\pm$ 0.017 0.49 $\pm$ 0.00
Sigmoid 0.003 0.01 - 0.76 $\pm$ 0.03 0.074 $\pm$ 0.008 0.49 $\pm$ 0.00
EasyMKL - - - 0.73 $\pm$ 0.01 0.084 $\pm$ 0.005 0.69 $\pm$ 0.01
UniformMK - - - 0.73 $\pm$ 0.01 0.084 $\pm$ 0.002 0.49 $\pm$ 0.00
Original - - - 1.56 0.000 $\geq 0.99$
Table 4: Evaluation of the neutralized pre-image representations. We show the WEAT test’s statistics and $p$-value, as well as the gender prediction accuracy of a kernel classifier of the same type as the one applied during neutralization.
A.5 Closest Neighbors
In Table 5, we show the closest neighbors to randomly-sampled word representations before and after gender erasure under the polynomial kernel. The results for other kernels are qualitatively similar.
Word Neighbors before Neighbors after
spiritual faith, religious, healing emotional, religious, healing
lesson learn, teach, lessons teaching, teach, lessons
faces faced, facing, face faced, facing, face
forget know, let, remember know, let, remember
converter ipod, conversion, convert ipod, conversion, convert
clean keep, wash, cleaning keep, wash, cleaning
formal elegant, dress, appropriate elegant, appropriate, dress
identity identify, context, identification context, identify, identification
other these, those, many these, those, many
licensed registered, certified, license registered, certified, license
ratings reviews, rated, rating reviews, rated, rating
properly proper, effectively, correctly effectively, proper, correctly
build create, built, building built, create, building
solutions systems, technologies, solution services, technologies, solution
afghanistan troops, pakistan, iraq troops, pakistan, iraq
wallpaper desktop, pictures, picture desktop, pictures, picture
sound audio, noise, sounds audio, noise, sounds
gender sexual, male, age male, differences, age
boat cruise, ship, fishing cruise, ship, fishing
downtown portland, city, neighborhood portland, neighborhood, city
lawyers attorney, lawyer, attorneys attorney, lawyer, attorneys
smart how, easy, intelligent wise, easy, intelligent
spending budget, spent, spend budget, spent, spend
contest winners, winner, competition winners, winner, competition
want n’t, know, need n’t, know, need
advice guidance, suggestions, tips guidance, suggestions, tips
professionals managers, professional, experts managers, professional, experts
g d, b, f d, b, f
australian zealand, british, australia zealand, british, australia
na mo, o, da mo, o, da
Table 5: Closest neighbors to randomly-sampled words from GloVe vocabulary, for the original representations, and for the pre-images after our intervention.
A.6 WEAT Results
Here we report the results of the WEAT test for the career and family-related words (Table 6) and art and mathematics-related words (Table 7).
Kernel $\gamma$ $\alpha$ d WEAT-$d$ $p$-value
Poly 0.05 0.8 2 0.72 $\pm$ 0.00 0.093 $\pm$ 0.002
Poly 0.05 1 2 0.72 $\pm$ 0.01 0.091 $\pm$ 0.003
Poly 0.05 1.2 2 0.72 $\pm$ 0.01 0.093 $\pm$ 0.004
Poly 0.05 0.8 3 0.73 $\pm$ 0.00 0.089 $\pm$ 0.001
Poly 0.05 1 3 0.74 $\pm$ 0.01 0.086 $\pm$ 0.004
Poly 0.05 1.2 3 0.73 $\pm$ 0.01 0.089 $\pm$ 0.002
Poly 0.1 0.8 2 0.72 $\pm$ 0.01 0.091 $\pm$ 0.005
Poly 0.1 1 2 0.73 $\pm$ 0.00 0.089 $\pm$ 0.001
Poly 0.1 1.2 2 0.72 $\pm$ 0.01 0.090 $\pm$ 0.003
Poly 0.1 0.8 3 0.72 $\pm$ 0.01 0.090 $\pm$ 0.003
Poly 0.1 1 3 0.72 $\pm$ 0.01 0.090 $\pm$ 0.003
Poly 0.1 1.2 3 0.72 $\pm$ 0.01 0.091 $\pm$ 0.002
Poly 0.15 0.8 2 0.71 $\pm$ 0.01 0.095 $\pm$ 0.004
Poly 0.15 1 2 0.74 $\pm$ 0.00 0.087 $\pm$ 0.001
Poly 0.15 1.2 2 0.72 $\pm$ 0.01 0.091 $\pm$ 0.002
Poly 0.15 0.8 3 0.72 $\pm$ 0.00 0.093 $\pm$ 0.001
Poly 0.15 1 3 0.72 $\pm$ 0.01 0.092 $\pm$ 0.002
Poly 0.15 1.2 3 0.73 $\pm$ 0.01 0.090 $\pm$ 0.005
RBF 0.1 - - 0.72 $\pm$ 0.01 0.090 $\pm$ 0.003
RBF 0.15 - - 0.73 $\pm$ 0.01 0.090 $\pm$ 0.005
RBF 0.2 - - 0.72 $\pm$ 0.01 0.091 $\pm$ 0.003
Laplace 0.1 - - 0.75 $\pm$ 0.05 0.083 $\pm$ 0.017
Laplace 0.15 - - 0.77 $\pm$ 0.02 0.076 $\pm$ 0.007
Laplace 0.2 - - 0.70 $\pm$ 0.02 0.098 $\pm$ 0.008
Linear - - - 0.72 $\pm$ 0.02 0.090 $\pm$ 0.006
Sigmoid 0.005 0 - 0.73 $\pm$ 0.03 0.087 $\pm$ 0.010
Sigmoid 0.005 0.01 - 0.73 $\pm$ 0.02 0.087 $\pm$ 0.006
Sigmoid 0.003 0 - 0.76 $\pm$ 0.06 0.079 $\pm$ 0.019
Sigmoid 0.003 0.01 - 0.76 $\pm$ 0.07 0.080 $\pm$ 0.020
EasyMKL - - - 0.72 $\pm$ 0.02 0.091 $\pm$ 0.005
UniformMK - - - 0.72 $\pm$ 0.00 0.092 $\pm$ 0.002
Original - - - 1.69 0.000
Table 6: Word association bias test (WEAT) for career and family-related terms
Kernel $\gamma$ $\alpha$ d WEAT-d $p$-value
Poly 0.05 0.8 2 0.78 $\pm$ 0.00 0.068 $\pm$ 0.001
Poly 0.05 1 2 0.78 $\pm$ 0.01 0.067 $\pm$ 0.002
Poly 0.05 1.2 2 0.78 $\pm$ 0.00 0.067 $\pm$ 0.001
Poly 0.05 0.8 3 0.78 $\pm$ 0.01 0.066 $\pm$ 0.002
Poly 0.05 1 3 0.78 $\pm$ 0.01 0.066 $\pm$ 0.002
Poly 0.05 1.2 3 0.78 $\pm$ 0.00 0.066 $\pm$ 0.001
Poly 0.1 0.8 2 0.78 $\pm$ 0.00 0.068 $\pm$ 0.001
Poly 0.1 1 2 0.78 $\pm$ 0.01 0.066 $\pm$ 0.002
Poly 0.1 1.2 2 0.77 $\pm$ 0.01 0.069 $\pm$ 0.004
Poly 0.1 0.8 3 0.78 $\pm$ 0.00 0.067 $\pm$ 0.001
Poly 0.1 1 3 0.77 $\pm$ 0.01 0.070 $\pm$ 0.001
Poly 0.1 1.2 3 0.78 $\pm$ 0.01 0.068 $\pm$ 0.002
Poly 0.15 0.8 2 0.78 $\pm$ 0.01 0.068 $\pm$ 0.003
Poly 0.15 1 2 0.78 $\pm$ 0.01 0.068 $\pm$ 0.003
Poly 0.15 1.2 2 0.78 $\pm$ 0.00 0.068 $\pm$ 0.001
Poly 0.15 0.8 3 0.78 $\pm$ 0.01 0.067 $\pm$ 0.002
Poly 0.15 1 3 0.78 $\pm$ 0.01 0.067 $\pm$ 0.002
Poly 0.15 1.2 3 0.77 $\pm$ 0.01 0.069 $\pm$ 0.002
RBF 0.1 - - 0.79 $\pm$ 0.01 0.066 $\pm$ 0.003
RBF 0.15 - - 0.78 $\pm$ 0.01 0.067 $\pm$ 0.002
RBF 0.2 - - 0.78 $\pm$ 0.01 0.067 $\pm$ 0.002
Laplace 0.1 - - 0.80 $\pm$ 0.04 0.064 $\pm$ 0.012
Laplace 0.15 - - 0.81 $\pm$ 0.03 0.061 $\pm$ 0.009
Laplace 0.2 - - 0.77 $\pm$ 0.04 0.070 $\pm$ 0.013
linear - - - 0.79 $\pm$ 0.01 0.066 $\pm$ 0.002
Sigmoid 0.005 0 - 0.82 $\pm$ 0.04 0.057 $\pm$ 0.010
Sigmoid 0.005 0.01 - 0.77 $\pm$ 0.05 0.070 $\pm$ 0.012
Sigmoid 0.003 0 - 0.76 $\pm$ 0.03 0.073 $\pm$ 0.009
Sigmoid 0.003 0.01 - 0.79 $\pm$ 0.03 0.066 $\pm$ 0.009
EasyMKL - - - 0.78 $\pm$ 0.00 0.066 $\pm$ 0.001
UniformMK - - - 0.78 $\pm$ 0.01 0.069 $\pm$ 0.002
Original - - - 1.56 0.000
Table 7: Word association bias test (WEAT) for art and mathematics-related terms
A.7 Transfer Results
UniformMK EasyMKL RBF Poly Laplace Sigmoid Linear MLP
poly, $\gamma=$0.05, d= 2, $\alpha=$0.8 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.01 0.49 $\pm$ 0.00 0.61 $\pm$ 0.06 0.97 $\pm$ 0.00
poly, $\gamma=$0.05, d= 2, $\alpha=$1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.01 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.52 $\pm$ 0.01 0.97 $\pm$ 0.00
poly, $\gamma=$0.05, d= 2, $\alpha=$1.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.53 $\pm$ 0.02 0.97 $\pm$ 0.00
poly, $\gamma=$0.05, d= 3, $\alpha=$0.8 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.01 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.49 $\pm$ 0.00 0.54 $\pm$ 0.03 0.97 $\pm$ 0.00
poly, $\gamma=$0.05, d= 3, $\alpha=$1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.01 0.49 $\pm$ 0.00 0.59 $\pm$ 0.05 0.97 $\pm$ 0.00
poly, $\gamma=$0.05, d= 3, $\alpha=$1.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.53 $\pm$ 0.01 0.97 $\pm$ 0.00
poly, $\gamma=$0.1, d= 2, $\alpha=$0.8 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.49 $\pm$ 0.00 0.54 $\pm$ 0.04 0.97 $\pm$ 0.00
poly, $\gamma=$0.1, d= 2, $\alpha=$1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.49 $\pm$ 0.00 0.53 $\pm$ 0.02 0.97 $\pm$ 0.00
poly, $\gamma=$0.1, d= 2, $\alpha=$1.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.54 $\pm$ 0.01 0.97 $\pm$ 0.00
poly, $\gamma=$0.1, d= 3, $\alpha=$0.8 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.54 $\pm$ 0.00 0.97 $\pm$ 0.00
poly, $\gamma=$0.1, d= 3, $\alpha=$1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.56 $\pm$ 0.02 0.97 $\pm$ 0.00
poly, $\gamma=$0.1, d= 3, $\alpha=$1.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.54 $\pm$ 0.03 0.97 $\pm$ 0.00
poly, $\gamma=$0.15, d= 2, $\alpha=$0.8 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.57 $\pm$ 0.02 0.97 $\pm$ 0.00
poly, $\gamma=$0.15, d= 2, $\alpha=$1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.54 $\pm$ 0.01 0.97 $\pm$ 0.00
poly, $\gamma=$0.15, d= 2, $\alpha=$1.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.54 $\pm$ 0.01 0.97 $\pm$ 0.00
poly, $\gamma=$0.15, d= 3, $\alpha=$0.8 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.49 $\pm$ 0.00 0.58 $\pm$ 0.03 0.97 $\pm$ 0.00
poly, $\gamma=$0.15, d= 3, $\alpha=$1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.57 $\pm$ 0.01 0.97 $\pm$ 0.00
poly, $\gamma=$0.15, d= 3, $\alpha=$1.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.56 $\pm$ 0.01 0.97 $\pm$ 0.00
rbf, $\gamma=$0.1 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.58 $\pm$ 0.04 0.97 $\pm$ 0.00
rbf, $\gamma=$0.15 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.60 $\pm$ 0.03 0.97 $\pm$ 0.00
rbf, $\gamma=$0.2 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.49 $\pm$ 0.00 0.60 $\pm$ 0.04 0.97 $\pm$ 0.00
laplace, $\gamma=$0.1 0.60 $\pm$ 0.06 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.01 0.49 $\pm$ 0.00 0.61 $\pm$ 0.03 0.97 $\pm$ 0.00
laplace, $\gamma=$0.15 0.56 $\pm$ 0.04 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.61 $\pm$ 0.01 0.97 $\pm$ 0.00
laplace, $\gamma=$0.2 0.59 $\pm$ 0.06 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.62 $\pm$ 0.02 0.97 $\pm$ 0.00
linear 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.93 $\pm$ 0.00 0.49 $\pm$ 0.00 0.54 $\pm$ 0.02 0.97 $\pm$ 0.00
sigmoid, $\gamma=$0.005, $\alpha=$0 0.64 $\pm$ 0.05 0.97 $\pm$ 0.00 0.93 $\pm$ 0.01 0.98 $\pm$ 0.00 0.90 $\pm$ 0.01 0.49 $\pm$ 0.00 0.65 $\pm$ 0.05 0.97 $\pm$ 0.00
sigmoid, $\gamma=$0.005, $\alpha=$0.01 0.63 $\pm$ 0.02 0.97 $\pm$ 0.00 0.94 $\pm$ 0.00 0.98 $\pm$ 0.00 0.90 $\pm$ 0.01 0.49 $\pm$ 0.00 0.65 $\pm$ 0.04 0.97 $\pm$ 0.00
sigmoid, $\gamma=$0.003, $\alpha=$0 0.63 $\pm$ 0.07 0.97 $\pm$ 0.00 0.92 $\pm$ 0.01 0.98 $\pm$ 0.00 0.89 $\pm$ 0.01 0.49 $\pm$ 0.00 0.66 $\pm$ 0.07 0.97 $\pm$ 0.00
sigmoid, $\gamma=$0.003, $\alpha=$0.01 0.65 $\pm$ 0.03 0.97 $\pm$ 0.00 0.92 $\pm$ 0.01 0.98 $\pm$ 0.00 0.89 $\pm$ 0.01 0.49 $\pm$ 0.00 0.65 $\pm$ 0.05 0.97 $\pm$ 0.00
EasyMKL 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.49 $\pm$ 0.00 0.57 $\pm$ 0.03 0.97 $\pm$ 0.00
UniformMK 0.49 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.01 0.49 $\pm$ 0.00 0.58 $\pm$ 0.08 0.97 $\pm$ 0.00
Table 8: Gender prediction from the neutralized pre-image representations using non-linear adversaries that differ from the neutralizing kernel.
In this appendix, we provide gender prediction accuracy, on the neutralized pre-image representations, with predictors that are different from those used in training (Experiment § 5.2).
Setup.
After projecting out the gender concept in kernel space, and computing the pre-image of the neutralized representations, we apply different non-linear kernels as well as an MLP to predict gender. We use the following parameters:
  • RBF : $\gamma=0.3$.
  • Poly : $\texttt{d}\ =3$, $\gamma=0.5$, $\alpha=0.3$.
  • Laplace : $\gamma=0.3$.
  • Sigmoid : $\alpha=0$, $\gamma=0.01$.
  • MLP : A network with a single 128-dimensional hidden layer with ReLU activations.
All classifiers were trained using sklearn.
Results.
The results are shown in Table 8. Rows denote the kernel that was applied for neutralization in Eq. 10, while columns denote the type of adversarial classifier applied on the final pre-image representations. Numbers denote accuracy in gender prediction.